Given the following utility function:

U(992)=9, +10 log2



Where, q1 and q2 are consumer goods and the budget constraint is given as.

U(992)=9, +10 log2


With p, and p the prices of the goods and the month the income. Find.

1. The Marshallian Demands for (q1 and 92.
2. The Indirect Utility Function, V (p1, p2, m)
3. The Hicksian Demands for q1 and q2.
4. The Expenditure Function, m (p1, p2, U)
Answer #1

ucan, an) = q + 10 log 92 and m. P. and P2 Ane (1) fog marshallion Demand functions, 9, and we max u= 9, + 10 log Q2 Sit m= bNow P,.qit bioait p2-92 = topi = m. m. prl Prod = m- lopi a = m-lopi = m-lopi - Pipi from lopi = 92. P2 - Q2 = lopi P2. q = mAns-3 Hicksian demand functions max. m= P, .91+ P2-92 st u= 9, + 10 log2 L = p.ait pr. q2t, pi- = 0 = = pi. 391 the = P2 - 1U = q + 100 bi p2 - =q, U - 100 PI p2 - Ub-loobi =q, where as Q2 = lobi p2 So ricksian demand 9 = functions UP2-loopi = U-loo